Circular No. 8819

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2007 E2 Terry Lovejoy, Thornlands, Queensland, Australia, reports his discov-

ery of a comet with a strong central condensation and a green extended 4′ coma, with a slight extension to the southwest, on sixteen 90-s CCD images obtained on Mar. 15 with a Canon 350D camera (+ 200-mm f/2.8 lens); the coma diameter was given as 5′ from his Mar. 16 frame (his positions below have estimated uncertainties of ± 15′′). At Lovejoy’s request, J. Drummond (Gisborne, New Zealand, 0.41-m reflector) confirmed the comet visually at mag 9.5 (moderately condensed, coma diameter 2′.6, no tail).

2007 UT α_{2000} δ_{2000} Mag. Observer
Mar. 15.727 20^{h}44^{m}05.^{s}5 −51^{o}14′14′′  10 Lovejoy
15.736 20 44 06.1 −51 13 22  _{”}
16.525 20 43 02.7 −50 49 35  9.5 Drummond
16.625 20 42 59.5 −50 45 39  _{”}
16.704 20 42 55.2 −50 43 14  9.3 Lovejoy

COMETS C/2006 X11 AND C/2006 Y3–Y9 (SOHO) Eight additional Kreutz sungrazing comets have been found on SOHO

website images (cf. IAUC 8815). All were stellar in appearance except C/2006 X11 (diffuse and small), C/2006 Y8 (slightly diffuse, tiny, mag ∼ 8), and C/2006 Y9 (“mostly stellar” in appearance). C/2006 X11, C/2006 Y3 (C3 images), C/2006 Y7, and C/2006 Y9 were very faint (mag ∼ 8.5). C/2006 Y3 was of mag ∼ 8 in C2 images. C/2006 Y4 was of mag ∼ 7.5–8. C/2006 Y5 reached mag 5.3 on Dec. 21.529 UT (at 10R_{⊙}) but showed no apparent tail. C/2006 Y6 reached mag 5.0 on Dec. 21.846 (at 11R_{⊙}) in C3 images, while C2 images show a hint of an extremely faint, thin tail.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 X11 Dec. 15.371 17^{h}33.^{m}8 −25^{o}05′ C2  RM 2007-D16
C/2006 Y3  17.863 17 47.8 −25 35 C3/2 SY 2007-D16
C/2006 Y4 20.038 17 55.6 −25 16 C2 RK 2007-D16
C/2006 Y5  20.929 18 00.8 −27 40 C3/2 TH 2007-D16
C/2006 Y6  21.238 18 02.3 −27 58 C3/2 BZ 2007-D17
C/2006 Y7  21.854 18 03.6 −25 20 C2  TH 2007-D17
C/2006 Y8  21.884 18 03.8 −25 10 C2  RK 2007-D17
C/2006 Y9  21.914 18 04.1 −25 15 C2  MA 2007-D17
2007 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green